What is difference between rotational kinetic energy and linear kinetic energy. Write down their formulas and compare them.

1
Expert's answer
2022-01-25T06:51:40-0500

Explanation


  • Kinetic energy is the energy associated with the motion of a body.
  • Linear kinetic energy describes the energy of an object that moves in linearly.
  • Whereas rotational kinetic energy is the energy, related to an object that spins about an axis.
  • The Linear kinetic energy is given by

"\qquad\qquad
\begin{aligned}
\small E_k&=\small \frac{1}{2}mv^2
\end{aligned}"

  • That for the rotational motion is

"\qquad\qquad
\begin{aligned}
\small E_{k-r}&=\small \frac{1}{2}I \omega^2
\end{aligned}"

  • Yet the 2 formulas are analogous and hence "\small I \implies m,\quad \omega\implies v" .